# A Researcher’s Guide to Empirical Risk Minimization

### v1

@@ 3.1 Step 1: the basic inequality (deterministic regret bound)
The basic inequality in Theorem 1 reduces the regret analysis to controlling the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$. The remainder of the argument is to bound this term with high probability and then translate the resulting bound into a rate for $R(\hat{f}_{n})-R(f_{0})$. We begin with a warm-up that builds intuition for the behavior of this term and previews the main proof techniques used in ERM analyses.
@@ 3.2 Warm-up: turning the basic inequality into rates
This section illustrates, at a heuristic level, how the basic inequality in Theorem 1 can be converted into a rate for the regret $R(\hat{f}_{n})-R(f_{0})$. The key term is the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$: once this quantity is controlled (in probability or with high probability), the bound translates directly into a regret rate.
It is helpful to separate the sources of randomness in the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$. There are two sources: (i) $P_{n}$ averages over an i.i.d. sample, and (ii) the function being averaged is random because it depends on the data through the ERM $\hat{f}_{n}$. Any ERM analysis must account for both effects. The first source alone can be handled using standard concentration inequalities for i.i.d. averages of the form $(P_{n}-P)h$ with $h$ fixed. The second source is the main complication: because $\hat{f}_{n}$ is data dependent, this fixed-function viewpoint is no longer sufficient, and one needs uniform control of the empirical process
@@
over the class $\mathcal{F}$. Indeed, $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ is obtained by evaluating this process at the data-dependent index $f=\hat{f}_{n}$. We therefore begin with a warm-up in the fixed-function setting, isolating source (i), to illustrate how regret bounds are derived. Handling source (ii) requires more sophisticated uniform (typically local) concentration tools, but once such bounds are available, the overall ERM analysis follows the same proof template with only minor modifications.
@@
As the following lemma shows, one can in fact obtain an exponential-tail (PAC-style) bound for the empirical-process fluctuation, and hence for the regret.
Lemma 2 (Bernstein bound for a fixed loss difference).
@@
Rearranging yields $\{R(f_{n})-R(f_{0})\}^{1/2}=O_{p}(n^{-1/2})$, and therefore $R(f_{n})-R(f_{0})=O_{p}(n^{-1})$.
@@
Much of the intuition from the deterministic thought experiment carries over to ERM, where the random minimizer $\hat{f}_{n}$ satisfies the basic inequality. The main complication is that Lemma 2 controls the empirical-process fluctuation only for a fixed (deterministic) function, so it does not apply directly to the data-dependent choice $\hat{f}_{n}$. Extending the argument therefore requires high-probability bounds that hold uniformly over $f\in\mathcal{F}$. As we will see, such uniformity can be costly: outside low-dimensional parametric models, it typically rules out the optimistic $O_{p}(n^{-1})$ regret rate, and slower rates are the norm.
@@
This removes the data dependence of $\hat{f}_{n}$, at the cost of controlling the worst-case value of the empirical process. When the class $\mathcal{F}$ is not too large (e.g., when it is Donsker), one can often apply a uniform LLN/CLT to show that the supremum is $O_{p}(n^{-1/2})$, and hence $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}=O_{p}(n^{-1/2})$ (Van Der Vaart and Wellner, 1996). This recovers the “slow rate” $R(\hat{f}_{n})-R(f_{0})=O_{p}(n^{-1/2})$ that we argued heuristically when treating $\hat{f}_{n}$ as deterministic. However, just as in the deterministic case, this approach does not exploit the localization of $\hat{f}_{n}$ around $f_{0}$, since it replaces $\hat{f}_{n}$ with a supremum over all $f\in\mathcal{F}$.
@@ 3.3 A three-step template for ERM rates
The template consists of three steps. First, derive a deterministic basic inequality, as we already did in Theorem 1. Second, obtain a uniform concentration bound for the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ appearing in the basic inequality. To obtain fast rates, this bound should ideally reflect the localization of $\hat{f}_{n}$ around $f_{0}$, as illustrated in the previous section. Third, combine the basic inequality with the high-probability bound to obtain regret bounds for the ERM.
@@ 4.1 A general high-probability regret theorem
$\mathfrak{R}_{n}(\mathcal{G},\delta):=E\left[\sup_{\begin{subarray}{c}f\in\mathcal{G}\\ \|f\|\leq\delta\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Z_{i})\right],$
where $\epsilon_{1},\ldots,\epsilon_{n}\in\{-1,1\}$ are i.i.d. Rademacher random variables, independent of $Z_{1},\ldots,Z_{n}$. Define the star hull of $\mathcal{G}$ by $\mathrm{star}(\mathcal{G}):=\{\,th:h\in\mathcal{G},\ t\in[0,1]\,\}$. Any convex class containing 0 is star-shaped, and hence equals its star hull. A useful feature of $\mathrm{star}(\mathcal{G})$, which we use repeatedly in our proofs, is that it ensures the scaling property: for all $t\in[0,1]$,
@@
and hence $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\mathcal{G}),\delta)/\delta$ is nonincreasing (Lemma 10).
@@
Larger classes typically have larger critical radii. Intuitively, $\delta_{n}(\mathcal{G})$ marks the scale above which the empirical $L^{2}(P_{n})$ norm and the population $L^{2}(P)$ norm are comparable on $\mathcal{G}$ (see Theorem 14.1 of Wainwright, 2019). Roughly, when $g\in\mathcal{G}$ satisfies $\|g\|\asymp\delta_{n}(\mathcal{G})$, its empirical norm $\|g\|_{n}$ is typically of the same order, namely $\delta_{n}(\mathcal{G})$. Below this scale, sampling noise can dominate: $\|g\|_{n}$ may be much larger than $\|g\|$ purely by chance. For this reason, critical radii enter uniform high-probability bounds for empirical processes (see Theorem 10 in Appendix B). Critical radii are known for many function classes; common examples are given in Table 1. In the next section, we develop general tools for upper bounding the critical radius via metric entropy bounds on localized Rademacher complexity.
Table 1: Examples of critical radii for common function classes.
Function class $\mathcal{G}$ Critical radius $\delta_{n}$ Reference
To bound the regret, we first establish a uniform high-probability bound of the form E1 controlling the local concentration of the empirical process $\{(P_{n}-P)g:g\in\mathcal{F}_{\ell}\}$. The resulting bound is governed by the critical radius of the centered loss-difference class
@@
The bound and its proof follow Bartlett et al. (2005), leveraging that the map $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\mathcal{G}),\delta)/\delta$ is nonincreasing. Alternative local concentration inequalities include Theorem 14.20 of Wainwright (2019) (see also Appendix K of Foster and Syrgkanis, 2023).
@@
Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$. Define $B:=\sup_{f\in\mathcal{F}}\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|_{\infty}$ and $\sigma_{f}^{2}:=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f)\}$ for each $f\in\mathcal{F}$. For all $\eta\in(0,1)$, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\sigma_{f}\delta_{n}+\delta_{n}^{2}+\sigma_{f}(B\vee 1)\sqrt{\frac{\log(1/\eta)}{n}}+(B\vee 1)^{2}\frac{\log(1/\eta)}{n}\right].$
@@
Our main result is the following PAC-style regret bound. Its proof follows the template of Section 3.3, using that Condition E1 holds with $\varepsilon(n,\sigma_{f},\eta)$ given by the right-hand side of Theorem 2. We assume the following high-level Bernstein condition; sufficient conditions are provided in Lemma 3.
@@ 1. A1)
When working with the $L^{2}$ error and a stronger Bernstein condition, applying the second part of Theorem 2 yields a slightly cleaner bound than Theorem 3, in which the critical radius may be taken with respect to the uncentered loss-difference class $\mathrm{star}(\mathcal{F}_{\ell})$ at any sample size.
@@ 1. A2)
Theorem 3 reduces bounding the regret $R(\hat{f}_{n})-R(f_{0})$ to controlling the localized Rademacher complexity—and the corresponding critical radius—of the star hull of the centered loss-difference class $\overline{\mathcal{F}}_{\ell}$. As a result, the regret converges to zero at rate $O_{p}(\delta_{n}^{2})$, up to a rapidly vanishing and typically negligible $O_{p}(n^{-1})$ term. For strongly convex losses over convex classes, Lemma 1 further yields the $L^{2}$ error bound $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})+O_{p}(n^{-1/2})$. Moreover, if $R(\hat{f}_{n})-R(f_{0})=o_{p}(1)$, the second part of Theorem 3 allows one to work with the uncentered loss-difference class: it suffices to choose $\delta_{n}$ so that $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$. With this choice, the same high-probability bound holds for $n$ large enough, yielding $R(\hat{f}_{n})-R(f_{0})=O_{p}(\delta_{n}^{2})+O_{p}(n^{-1})$. When working directly with $L^{2}$ error, Theorem 4 shows that $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})+O_{p}(n^{-1/2})$ under the same choice of $\delta_{n}$, with the corresponding high-probability bound valid at any sample size.

### v2

@@ 3.1 Step 1: the basic inequality (deterministic regret bound)
The basic inequality in Theorem 1 reduces the regret analysis to controlling the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$. The remainder of the argument is to bound this term and translate the resulting bound into a rate for $R(\hat{f}_{n})-R(f_{0})$.
@@ 3.2 Warm-up: turning the basic inequality into rates
This section illustrates, at a heuristic level, how the basic inequality yields regret rates. The goal is to build intuition for the empirical-process fluctuation term $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ and to preview the main proof techniques used in ERM analyses.
It is helpful to distinguish the two sources of randomness in the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$. First, $P_{n}$ averages over an i.i.d. sample. Second, the function being averaged is itself random because it depends on the data through the ERM $\hat{f}_{n}$. Any ERM analysis must account for both effects. The first source alone can be handled using standard concentration inequalities for i.i.d. averages of the form $(P_{n}-P)h$ with $h$ fixed. The second source is the main complication: because $\hat{f}_{n}$ is data dependent, this fixed-function viewpoint is no longer sufficient, and one instead needs uniform control of the empirical process
@@
over the class $\mathcal{F}$. Indeed, $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ is obtained by evaluating this process at the data-dependent index $f=\hat{f}_{n}$. We therefore begin with a warm-up in the fixed-function setting, isolating source (i), to illustrate how regret bounds are derived. Handling source (ii) requires more sophisticated uniform, typically local, concentration tools, but once such bounds are available, the overall ERM analysis follows the same proof template with only minor modifications.
@@
As the following lemma shows, one can in fact obtain an exponential-tail, PAC-style bound for the empirical-process fluctuation, and hence for the regret; see Sridharan (2002) for an introduction to concentration inequalities.
Lemma 2 (Bernstein-type point-wise local concentration bound).
@@
Rearranging yields $\{R(f_{n})-R(f_{0})\}^{1/2}=O_{p}(n^{-1/2})$, and therefore $R(f_{n})-R(f_{0})=O_{p}(n^{-1})$. This $n^{-1}$ rate is generally unattainable in nonparametric classes, because uniform control of the empirical process introduces additional complexity terms.
@@
Much of the intuition from the deterministic thought experiment carries over to ERM, where the random minimizer $\hat{f}_{n}$ satisfies the basic inequality. The main complication is that Lemma 2 controls the empirical-process fluctuation only for a fixed (deterministic) function, and therefore does not apply directly to the data-dependent choice $\hat{f}_{n}$. Extending the argument requires high-probability bounds that hold uniformly over $f\in\mathcal{F}$. As we will see, such uniformity is typically costly: outside low-dimensional parametric models, it rules out the optimistic $O_{p}(n^{-1})$ regret rate, and slower rates are the norm.
@@
This removes the data dependence of $\hat{f}_{n}$, at the cost of controlling the worst-case fluctuation of the empirical process over the entire class. When $\mathcal{F}$ is not too large (for example, when it is Donsker), one can often apply a uniform law of large numbers or central limit theorem (Van Der Vaart and Wellner, 1996) to show that this supremum is $O_{p}(n^{-1/2})$, and hence
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}=O_{p}(n^{-1/2}).$
This recovers the “slow rate” $R(\hat{f}_{n})-R(f_{0})=O_{p}(n^{-1/2})$ that we argued heuristically in the deterministic case. However, as in the deterministic setting, this approach does not exploit the localization of $\hat{f}_{n}$ around $f_{0}$, since it replaces $\hat{f}_{n}$ by a supremum over all $f\in\mathcal{F}$.
@@ 3.3 A three-step template for ERM rates
The template consists of three steps. First, derive a deterministic basic inequality, as we already did in Theorem 1. Second, obtain a high probability bound for the empirical-process fluctuation $(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,\hat{f}_{n})\}$ appearing in the basic inequality. To obtain fast rates, this bound should ideally reflect the localization of $\hat{f}_{n}$ around $f_{0}$, as illustrated in the previous section. Third, combine the basic inequality with the high-probability bound to obtain regret bounds for the ERM.
@@ 4.1 A general high-probability regret theorem
$\mathfrak{R}_{n}(\mathcal{G},\delta):=\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in\mathcal{G}\\ \|f\|\leq\delta\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(Z_{i})\right],$
where $\epsilon_{1},\ldots,\epsilon_{n}\in\{-1,1\}$ are i.i.d. Rademacher random variables, independent of $Z_{1},\ldots,Z_{n}$. Intuitively, this quantity measures how well functions in $\mathcal{G}$ can align their values $f(Z_{i})$ with random signs $\epsilon_{i}$. If the supremum is large, then $\mathcal{G}$ is flexible enough to fit noise in the sample and is therefore more prone to overfitting; if it is small, then $\mathcal{G}$ is more constrained and should generalize better. If the supremum is effectively taken over a single function $g$ with $\|g\|\leq\delta$, then typically $\mathfrak{R}_{n}(\mathcal{G},\delta)\asymp\delta n^{-1/2}$. The same scaling arises for low-dimensional parametric classes. For many nonparametric classes, by contrast, the dependence on $\delta$ is slower, often of the form $\delta^{\alpha}n^{-1/2}$ for some $\alpha\in(0,1)$; see Table 1.
Generalization bounds, such as those in Wainwright (2019), often involve star-shaped classes because localized Rademacher complexity satisfies a useful scaling property on such classes. We define the star hull of $\mathcal{G}$ as
$\mathrm{star}(\mathcal{G}):=\{\,th:h\in\mathcal{G},\ t\in[0,1]\,\}.$
Any convex class containing 0 is star-shaped and therefore equals its star hull. A key feature of $\mathrm{star}(\mathcal{G})$, which we use repeatedly in our proofs, is that for all $t\in[0,1]$,
@@
and hence $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\mathcal{G}),\delta)/\delta$ is nonincreasing; see Lemma 10 in Appendix B. In words, shrinking functions in $\mathrm{star}(\mathcal{G})$ toward zero by a factor $t\in[0,1]$ reduces the local complexity at least proportionally.
@@
Larger classes typically have larger critical radii. Intuitively, $\delta_{n}(\mathcal{G})$ marks the scale at which the empirical $L^{2}(P_{n})$ norm and the population $L^{2}(P)$ norm begin to be comparable on $\mathcal{G}$; see Theorem 14.1 of Wainwright, 2019. Roughly, when $g\in\mathcal{G}$ satisfies $\|g\|\asymp\delta_{n}(\mathcal{G})$, its empirical norm $\|g\|_{n}$ is typically of the same order. Below this scale, sampling noise can dominate, so $\|g\|_{n}$ may be much larger than $\|g\|$ purely by chance. For this reason, critical radii enter uniform local concentration inequalities for empirical processes; see Theorem 10 in Appendix B. Localized complexities and critical radii are known for many function classes, and common examples are given in Table 1. In the next section, we develop general tools for upper bounding the critical radius via metric entropy bounds on localized Rademacher complexity.
Table 1: Examples of localized Rademacher complexity scalings $\mathfrak{R}_{n}(\mathcal{G},\delta)$ and critical radii $\delta_{n}$ for common function classes. The reported local complexities are upper bounds, and the displayed scaling may hold only for $\delta$ larger than the critical radius $\delta_{n}$. See Appendix A.1 for references.
Function class $\mathcal{G}$ Local scaling Critical radius $\delta_{n}$
Star hull of single function $\mathrm{star}(\{f\})$ $n^{-1/2}\delta$ $n^{-1/2}$
$s$-sparse linear predictors in $\mathbb{R}^{p}$ $n^{-1/2}\delta\sqrt{s\log(ep/s)}$ $\sqrt{s\log(ep/s)/n}$
VC-subgraph of dimension $V$ $n^{-1/2}\delta\sqrt{V\log(1/\delta)}$ $\sqrt{V\log(n/V)/n}$
Hölder/Sobolev with smoothness $s$ in dimension $d<2s$ $n^{-1/2}\delta^{1-d/(2s)}$ $n^{-s/(2s+d)}$
Bounded Hardy–Krause variation $n^{-1/2}\delta^{1/2}(\log(1/\delta))^{d-1}$ $n^{-1/3}(\log n)^{2(d-1)/3}$
RKHS with eigenvalue decay $\sigma_{j}\asymp j^{-2\alpha},\alpha>1/2$ $n^{-1/2}\delta^{1-1/(2\alpha)}$ $n^{-\alpha/(2\alpha+1)}$
To bound the regret, we first establish a uniform local concentration inequality for the empirical process $\{(P_{n}-P)g:g\in\mathcal{F}_{\ell}\}$ of the form in Condition E1. The resulting bound is governed by the critical radius of the centered loss-difference class
@@
In the next section, we study how to bound the associated critical radius. The result and its proof follow Bartlett et al. (2005) and Bartlett and Mendelson (2006), exploiting the fact that the map $\delta\mapsto\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta)/\delta$ is nonincreasing. Alternative uniform local concentration inequalities are given in Theorem 14.20 of Wainwright (2019); see also Appendix K of Foster and Syrgkanis (2023). Appendix B.5 provides a general template for deriving such bounds from local maximal inequalities.
@@
Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$. Suppose there exists $B\in[1,\infty)$ such that $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|_{\infty}\leq B$, and define $\sigma_{f}^{2}:=\operatorname{Var}\{\ell(Z,f_{0})-\ell(Z,f)\}$ for each $f\in\mathcal{F}$. For all $\eta\in(0,1)$, there exists a universal constant $C>0$ such that, with probability at least $1-\eta$, for every $f\in\mathcal{F}$,
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\sigma_{f}\delta_{n}+\delta_{n}^{2}+\sigma_{f}B\sqrt{\frac{\log(1/\eta)}{n}}+B^{2}\frac{\log(1/\eta)}{n}\right].$
@@
Our main result is the following PAC-style regret bound. Its proof follows the template of Section 3.3, using that E1 holds with $\varepsilon(n,\sigma_{f},\eta)$ given by the right-hand side of Theorem 2. We assume the following high-level Bernstein condition; sufficient conditions are provided in Lemma 3.
@@ 1. A1)
When working instead with the $L^{2}$ error and under a stronger Bernstein condition, applying the second part of Theorem 2 yields a slightly cleaner bound than Theorem 3. In that case, the critical radius may be taken with respect to the uncentered loss-difference class $\mathrm{star}(\mathcal{F}_{\ell})$ at any sample size.
@@ 1. A2)
Theorem 3 reduces the problem of bounding the regret $R(\hat{f}_{n})-R(f_{0})$ to controlling the localized Rademacher complexity—and the corresponding critical radius—of the star hull of the centered loss-difference class $\overline{\mathcal{F}}_{\ell}$. Consequently, the regret converges to zero at rate $O_{p}(\delta_{n}^{2})$, up to a typically second-order $O_{p}(n^{-1})$ term. For strongly convex losses over convex classes, Lemma 1 further yields the $L^{2}$ error bound $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})+O_{p}(n^{-1/2})$. When working directly with $L^{2}$ error, Theorem 4 shows that $\|\hat{f}_{n}-f_{0}\|=O_{p}(\delta_{n})+O_{p}(n^{-1/2})$, where $\delta_{n}$ satisfies $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$, with the critical radius now taken with respect to the uncentered loss-difference class. Alternatively, if $R(\hat{f}_{n})-R(f_{0})=o_{p}(1)$, the second part of Theorem 3 permits this choice of $\delta_{n}$ for all sufficiently large $n$, yielding $R(\hat{f}_{n})-R(f_{0})=O_{p}(\delta_{n}^{2})+O_{p}(n^{-1})$.
@@
Aside: eventual almost-sure bounds
In some applications, one is mainly interested in an asymptotic rate rather than a sharp finite-sample probability bound. The following corollaries yield events that hold eventually almost surely, meaning that there exists a (potentially random) $N<\infty$ such that, for all $n\geq N$, the events hold almost surely; this can simplify proofs. The idea is simple: take $\eta\asymp n^{-2}$ in the high-probability bound and apply the Borel–Cantelli lemma. This introduces only a $\log n$ factor.
Corollary 1 (Eventual almost-sure local concentration bound).
Let $\delta_{n}>0$ satisfy the critical radius condition $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}$. Suppose the conditions of Theorem 2 hold. Then there exists a universal constant $C>0$ such that, with probability one, for all sufficiently large $n$, every $f\in\mathcal{F}$ satisfies
$(P_{n}-P)\{\ell(\cdot,f_{0})-\ell(\cdot,f)\}\leq C\left[\sigma_{f}\delta_{n}+\delta_{n}^{2}+\sigma_{f}B\sqrt{\frac{\log n}{n}}+B^{2}\frac{\log n}{n}\right].$
The same conclusion holds under the alternative critical-radius condition $\mathfrak{R}_{n}(\mathrm{star}(\mathcal{F}_{\ell}),\delta_{n})\leq\delta_{n}^{2}$, with $\sigma_{f}$ replaced by $\|\ell(\cdot,f_{0})-\ell(\cdot,f)\|$.
Corollary 2 (Eventual almost-sure regret bound).
Assume Condition A1. Suppose $\sup_{f\in\mathcal{F}}\|\ell(\cdot,f)\|_{\infty}\leq M$ for some $M\in[1,\infty)$, and let $\delta_{n}>0$ satisfy $\mathfrak{R}_{n}(\mathrm{star}(\overline{\mathcal{F}}_{\ell}),\delta_{n})\lesssim\delta_{n}^{2}.$ Then there exists a universal constant $C\in(0,\infty)$ such that, with probability one, for all sufficiently large $n$,
$R(\hat{f}_{n})-R(f_{0})\leq C\,c_{\mathrm{Bern}}\Biggl[\delta_{n}^{2}+M^{2}\frac{\log n}{n}\Biggr].$